# Locally Optimal Fixed-Budget Best Arm Identification in Two-Armed Gaussian Bandits with Unknown Variances

### 20 Dec 2023 v1

@@ 4 Misidentification Probability and Asymptotic Optimality
=$=\Big{\{}\lambda\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]+\frac{\lambda^{2}}{T}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}/2!|\mathcal{F}_{t-1}\right]+O\left(\lambda^{3}\right)\Big{\}}-\frac{1}{2}\left\{\lambda\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]+O\left(\lambda^{2}\right)\right\}^{2}$
$=\frac{\lambda^{2}}{T}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}/2!|\mathcal{F}_{t-1}\right]+O\left(\lambda^{3}\right).$
Here, we used $\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]=0$. Thus, the (5) holds.
@@
For any $P^{*}\in\mathcal{P}^{\mathrm{G}}$, any $\epsilon>0$, there exists $t(\epsilon)>0$ such that for all $T>t(\epsilon)$, we have
@@
Our proof refers to the proof of Lemma 10 in Hadad et al. (2021).
@@
$=\mathbb{E}_{P^{*}}\left[\frac{\sigma^{2}_{1}+(\mu_{1}-\widehat{\mu}_{1,t})^{2}}{\widehat{w}_{1,t}}|\mathcal{F}_{t-1}\right]+\mathbb{E}_{P^{*}}\left[\frac{\sigma^{2}_{2}+(\mu_{2}-\widehat{\mu}_{2,t})^{2}}{\widehat{w}_{2,t}}|\mathcal{F}_{t-1}\right]-\mathbb{E}_{P^{*}}\left[\left(\big{(}\widehat{\mu}_{1,t}-\widehat{\mu}_{2,t}\big{)}-\big{(}\mu_{1}-\mu_{2}\big{)}\right)^{2}|\mathcal{F}_{t-1}\right].$
We assumed that $(\mu_{1},\mu_{2},\widehat{\mu}_{1,t},\widehat{\mu}_{2.t},\widehat{w}_{1,t},\widehat{w}_{2,t})$ are all bounded random variables. Let $C$ be a constant independent of $T$ such that $|u_{t}|<C$ for all $t\in\mathbb{N}$.
@@
$\mathbb{P}_{P^{*}}\left(\sum^{T}_{t=1}\Psi_{t}\leq Tv\right)\leq\mathbb{E}_{P^{*}}\left[\exp\left(-\frac{T\lambda^{2}}{2}+\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}+TO\left(\lambda^{3}\right)\right)\right].$
@@
$\mathbb{E}_{P^{*}}\left[\exp\left(-\frac{T\lambda^{2}}{2}+\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}+TO\left(\lambda^{3}\right)\right)\right]$
$=\exp\left(-\frac{T\lambda^{2}}{2}+TO\left(\lambda^{3}\right)\right)\mathbb{E}_{P^{*}}\left[\exp\left(\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}\right)\right]$
$\leq\exp\left(-\frac{T\lambda^{2}}{2}+TO\left(\lambda^{3}\right)\right)\exp\left(\frac{\lambda^{2}}{2}T\epsilon\right)=\exp\left(-\frac{T\lambda^{2}}{2}+TO\left(\lambda^{3}\right)+\frac{\lambda^{2}}{2}T\epsilon\right).$
@@
For any $\epsilon>0$, there exists $t(\epsilon)>0$ such that for all $T>t(\epsilon)$, we obtain
$-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{\mu}^{\mathrm{AIPW}}_{1,T}\leq\widehat{\mu}^{\mathrm{AIPW}}_{2,T}\right)\geq-\left\{-\frac{\lambda^{2}}{2}+O\left(\frac{\lambda^{3}}{\sqrt{T}}\right)+\frac{\lambda^{2}}{2}\epsilon\right\}=\frac{\lambda^{2}}{2}-O\left(\lambda^{3}\right)-\frac{\lambda^{2}}{2}\epsilon,$
@@
$-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{\mu}^{\mathrm{AIPW},a^{*}(P)}_{T}\leq\widehat{\mu}^{\mathrm{AIPW},a}_{T}\right)\geq\frac{\Delta^{2}}{2V}-O\left(\frac{\Delta^{3}}{V^{3/2}}\right)-\frac{\epsilon\Delta^{2}}{2V},$
as $\Delta\to 0$. Letting $\Delta\to 0$ and $T\to\infty$, and then letting $\epsilon\to 0$ independently of $T$ and $\Delta$, then $T\to\infty$, we have

### 17 Mar 2024 v2

@@ 4 Misidentification Probability and Asymptotic Optimality
=$=\Big{\{}\lambda\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]+\frac{\lambda^{2}}{T}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}/2!|\mathcal{F}_{t-1}\right]+o\left(\lambda^{2}\right)\Big{\}}-\frac{1}{2}\left\{\lambda\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]+o\left(\lambda\right)\right\}^{2}$
$=\frac{\lambda^{2}}{T}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}/2!|\mathcal{F}_{t-1}\right]+o\left(\lambda^{2}\right)$
as $\lambda\to 0$. Here, we used $\mathbb{E}_{P^{*}}\left[\Psi_{t}|\mathcal{F}_{t-1}\right]=0$. Thus, the (4) holds.
@@
For any $P^{*}\in\mathcal{P}^{\mathrm{G}}$ and any $\epsilon>0$, there exists $t(\epsilon)>0$ such that for all $T>t(\epsilon)$, we have
@@
This result is a variant of the Cesàro lemma for a case with almost sure convergence. For completeness, we show the proof, which is based on the proof of Lemma 10 in Hadad et al. (2021).
@@
$=\mathbb{E}_{P^{*}}\left[\frac{\sigma^{2}_{1}+(\mu^{*}_{1}-\widehat{\mu}_{1,t})^{2}}{\widehat{w}_{1,t}}|\mathcal{F}_{t-1}\right]+\mathbb{E}_{P^{*}}\left[\frac{\sigma^{2}_{2}+(\mu^{*}_{2}-\widehat{\mu}_{2,t})^{2}}{\widehat{w}_{2,t}}|\mathcal{F}_{t-1}\right]-\mathbb{E}_{P^{*}}\left[\left(\big{(}\widehat{\mu}_{1,t}-\widehat{\mu}_{2,t}\big{)}-\big{(}\mu^{*}_{1}-\mu^{*}_{2}\big{)}\right)^{2}|\mathcal{F}_{t-1}\right].$
We assumed that $(\mu^{*}_{1},\mu^{*}_{2},\widehat{\mu}_{1,t},\widehat{\mu}_{2.t},\widehat{w}_{1,t},\widehat{w}_{2,t})$ are all bounded random variables. Let $C$ be a constant independent of $T$ such that $|u_{t}|<C$ for all $t\in\mathbb{N}$.
@@
$\mathbb{P}_{P^{*}}\left(\sum^{T}_{t=1}\Psi_{t}\leq Tv\right)\leq\mathbb{E}_{P^{*}}\left[\exp\left(-\frac{T\lambda^{2}}{2}+\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}+To\left(\lambda^{2}\right)\right)\right].$
@@
$\mathbb{E}_{P^{*}}\left[\exp\left(-\frac{T\lambda^{2}}{2}+\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}+To\left(\lambda^{2}\right)\right)\right]$
$=\exp\left(-\frac{T\lambda^{2}}{2}+To\left(\lambda^{2}\right)\right)\mathbb{E}_{P^{*}}\left[\exp\left(\frac{\lambda^{2}}{2}\left\{\sum^{T}_{t=1}\mathbb{E}_{P^{*}}\left[\Psi^{2}_{t}|\mathcal{F}_{t-1}\right]-1\right\}\right)\right]$
$\leq\exp\left(-\frac{T\lambda^{2}}{2}+To\left(\lambda^{2}\right)\right)\exp\left(\frac{\lambda^{2}}{2}T\epsilon\right)=\exp\left(-\frac{T\lambda^{2}}{2}+To\left(\lambda^{2}\right)+\frac{\lambda^{2}}{2}T\epsilon\right).$
@@
For any $\epsilon>0$, there exists $t(\epsilon)>0$ such that for all $T>t(\epsilon)$, we obtainz
$-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{\mu}^{\mathrm{AIPW}}_{1,T}\leq\widehat{\mu}^{\mathrm{AIPW}}_{2,T}\right)\geq-\left\{-\frac{\lambda^{2}}{2}+o\left(\lambda^{2}\right)+\frac{\lambda^{2}}{2}\epsilon\right\}=\frac{\lambda^{2}}{2}-o\left(\lambda^{2}\right)-\frac{\lambda^{2}}{2}\epsilon,$
@@
$-\frac{1}{T}\log\mathbb{P}_{P^{*}}\left(\widehat{\mu}^{\mathrm{AIPW},a^{\star}(P)}_{T}\leq\widehat{\mu}^{\mathrm{AIPW},a}_{T}\right)\geq\frac{\Delta^{2}}{2V}-o\left(\frac{\Delta^{2}}{V}\right)-\frac{\epsilon\Delta^{2}}{2V},$
as $\Delta\to 0$. By letting $\Delta\to 0$ and $T\to\infty$, and then letting $\epsilon\to 0$ independently of $T$ and $\Delta$, we have